A wall with a thermal conductivity of 0.50W/m·K is maintained at 40.0°C. The heat transfer through the wall is 250.0W. The wall surface area is 1.50m2 and its thickness is 1.00cm. Determine the temperature at the other surface.

1
Expert's answer
2022-04-11T12:05:32-0400

The Fourier's law says

"P=\kappa\frac{\Delta T}{\Delta x}A"

Hence, the temperature difference

"\Delta T=\frac{P\Delta x}{\kappa A}"

"\Delta T=\frac{250.0*0.01}{0.50*1.50}=3.33^\circ\rm C"

The temperature at the other surface

"T_2=T_1\pm\Delta T=40.0\pm3.33\\
=43.3^\circ\rm C \: \rm or\; 36.7^\circ\rm C"